# An observational signature for extremal black holes

### v2

@@ V.2 Numerical Results
$s^{K}_{\mathcal{I}^{+}}(\vartheta)=-\frac{1}{64M^{2}}\cdot\lim_{\tau\rightarrow\infty}\left[4M\tau^{2}\cdot\Psi(\tau,\vartheta)\\ -\frac{\tau}{\log(\tau/M)}\cdot\left(\lim_{\tau\rightarrow\infty}\tau^{2}\cdot\Psi(\tau,\vartheta)\right)+\frac{\tau^{3}}{\log(\tau/M)}\Psi(\tau,\vartheta)\right].$ (18)
This redefinition preserves the property that $s^{K}_{\mathcal{I}^{+}}=H[\psi]\neq 0$ for EK but $s^{K}_{\mathcal{I}^{+}}=0$ for sub-extremal Kerr.
Figure 4: Best-fit line demonstrating a linear relationship between the horizon constant $H$ and the asymptotically extracted $s^{K}_{\mathcal{I}^{+}}(\vartheta)$ with ${\theta^{1,2}}=\pi/2$ for extremal Kerr black holes. The slope of the best fit-line is $1.0028\pm 0.13$. Each data point is labelled by the initial data used for that computation i.e. the location and width of the Gaussian pulse. The Levenberg–Marquardt fitting errors are depicted by the vertical bars on each data point.
VI Conclusion
In this paper, we have investigated scalar perturbations of the Reissner-Nordström family and the Kerr family of black holes. Our numerical analysis has allowed us to derive a characteristic excitation for the radiation field at any given unit solid angle of future null infinity. The derived expression is non-zero only in the extremal case, where the black hole’s charge or angular momentum reaches its maximum value. This observation serves as an important observational signature for extremal black holes, distinguishing them from sub-extremal ones.
Furthermore, we have identified the extremal horizon instability as the source of this excitation in the radiation field. Remarkably, we have demonstrated that the magnitude of the excitation in the radiation field precisely matches the conserved horizon charge of the black hole. This shows that the horizon charge is classically measurable by external observers, including far-away observers, thus falsifying the no-hair hypothesis in the case of extremal black holes.
@@ Appendix A Appendix: Initial data with no support on the horizon
In this appendix, we show numerical results that offer numerical confirmation that the proposed signature $s_{\mathcal{I}^{+}}$ works correctly with the initial data lacks support on the horizon for both ERN and EK cases. In Fig. 5 we show a plot of $s_{\mathcal{I}^{+}}$ for both ERN and EK for an initial Gaussian centered at $\rho=4.0$ with width $0.22$ and truncated outside the window $[1.0,8.0]$. As mentioned previously, the horizon is located at $\rho=0.95$ and thus this data has no support on the horizon. It is clear that the signature $s_{\mathcal{I}^{+}}(\vartheta)$ asymptotes to 00 as we evolve the scalar field. The computational methodology for this Fig. 5 is the same as that used for Fig. 1.
Figure 5: Time evolution of $[4M/C_{1}(\vartheta)]\times s_{\mathcal{I}^{+}}(\vartheta)$ for ERN and EK for initial data with no support on the horizon. Note that $s_{\mathcal{I}^{+}}(\vartheta)$ asymptotes to 00 as we evolve the scalar field at $\mathcal{I}^{+}$.

### v3

@@ VI.2 Numerical Results
$s^{\color[rgb]{0,0,0}Kerr}_{\mathcal{I}^{+}}(\vartheta)=-\frac{1}{64M^{2}}\cdot\lim_{\tau\rightarrow\infty}\left[4M\tau^{2}\cdot\Psi(\tau,\vartheta)\\ -\frac{\tau}{\log(\tau/M)}\cdot\left(\lim_{\tau\rightarrow\infty}\tau^{2}\cdot\Psi(\tau,\vartheta)\right)+\frac{\tau^{3}}{\log(\tau/M)}\Psi(\tau,\vartheta)\right].$ (18)
This redefinition preserves the property that $s^{\color[rgb]{0,0,0}Kerr}_{\mathcal{I}^{+}}=H[\psi]\neq 0$ for EK but $s^{\color[rgb]{0,0,0}Kerr}_{\mathcal{I}^{+}}=0$ for sub-extremal Kerr.
It is worth noting that the specific numerical values of $s_{\mathcal{I}^{+}}(\vartheta)$ we obtained for both RN and Kerr in the subextremal cases were on the $10^{-2}$ scale at late times. We attribute those values not being closer to zero due to numerical errors arising from our data analysis methodology. This suggests our signature would only be able to conclusively make a claim on extremality if the measured value of $s_{\mathcal{I}^{+}}(\vartheta)$ (and thus $H[\psi]$) is much larger than $10^{-3}$ . However, we do not view this as a serious limitation of our work, since our goal is to offer an “in principle” test for extremality and not be constrained by the limitation of our current numerical analysis.
Figure 4: Best-fit line demonstrating a linear relationship between the horizon constant $H$ and the asymptotically extracted $s^{\color[rgb]{0,0,0}Kerr}_{\mathcal{I}^{+}}(\vartheta)$ with ${\theta^{1,2}}=\pi/2$ for extremal Kerr black holes. The slope of the best fit-line is $1.0\pm 0.1$. Each data point is labelled by the initial data used for that computation i.e. the location and width of the Gaussian pulse. The Levenberg–Marquardt fitting errors are depicted by the vertical bars on each data point.
VII Conclusion (RN & Kerr)
In this paper, we have investigated scalar perturbations of the Reissner-Nordström family and the Kerr family of black holes. Our numerical analysis has allowed us to derive a characteristic excitation for the radiation field at any given fixed angle of future null infinity. The derived expression is non-zero only in the extremal case, where the black hole’s charge or angular momentum reaches its maximum value. This observation serves as an important observational signature for extremal black holes, distinguishing them from sub-extremal ones.
Furthermore, we have identified the extremal horizon instability as the source of this excitation in the radiation field. Remarkably, we have demonstrated that the magnitude of the excitation in the radiation field precisely matches the conserved horizon charge of the black hole. This demonstrates that the horizon charge is classically measurable by external observers, including far-away observers, potentially falsifying the no-hair hypothesis in the case of extremal black holes in the context of linearized black hole perturbation theory.
@@ Appendix A Appendix: Initial data with no support on the horizon
In this appendix, we show numerical results that offer numerical confirmation that the proposed signature $s_{\mathcal{I}^{+}}$ works correctly with the initial data lacks support on the horizon for both ERN and EK cases. In Fig. 5 we show a plot of $s_{\mathcal{I}^{+}}$ for both ERN and EK for an initial Gaussian centered at $\rho=4.0$ with width $0.22$ and truncated outside the window $[1.0,8.0]$. As mentioned previously, the horizon is located at $\rho=0.95$ and thus this data has no support on the horizon. It is clear that the signature $s_{\mathcal{I}^{+}}(\vartheta)$ asymptotes to 0 as we evolve the scalar field. The computational methodology for this Fig. 5 is the same as that used for Fig. 1.
Figure 5: Time evolution of $s_{\mathcal{I}^{+}}(\vartheta)$ for ERN and EK for initial data with no support on the horizon. Note that $s_{\mathcal{I}^{+}}(\vartheta)$ asymptotes to 0 as we evolve the scalar field at $\mathcal{I}^{+}$.